Sarah can paint a fence in 4 hours. Her father Joe can paint it in 2 hours. If they work together, how long will it take them?
Answer by ptaylor(2198)

Let x=amt of time it takes Sarah and her father working together to paint the fence
Then Sarah and her father working together paints at the rate of 1/x of the fence per hour
Sarah paints at the rate of 1/4 of the fence per hour
Her father paints at the rate of 1/2 of the fence per hour
Sooooo
1/4 + 1/2 = 1/x multiply each term by 4x
x + 2x=4
3x=4
x=4/3 hr= 1 hr 20 min
CK
(4/3)(1/4)+(4/3)(1/2) should equal 1 (1 fence, that is)
(1/3) +(2/3)=1
1=1
Hope this helps--- ptaylor